# FairGE: Fairness-Aware Graph Encoding in Incomplete Social Networks

### v1

@@ 1. Introduction
Graph Transformers (GTs) address common Graph Neural Networks (GNNs) challenges such as over-smoothing (Zhou et al., 2025), over-squash-ing (Gravina et al., 2025), and long-range dependencies (Chen et al., 2024b) by employing global attention to capture long-range node interactions (Wu et al., 2024; Zhang et al., 2025). Despite these advantages, GTs often overlook inherent biases in graph data, leading to discriminatory predictions towards sensitive subgroups (e.g., gender and race) (Luo et al., 2024b). This issue is particularly critical in social networks, where fairness directly affects user trust and system reliability (Zhong et al., 2024; Bernard and Balog, 2025). Here, algorithmic fairness requires that model predictions remain unbiased with respect to sensitive attributes (e.g., gender, race, region, and other potential factors), ensuring equitable treatment of diverse subgroups (Dong et al., 2023). The challenge is further compounded by the frequent absence of sensitive attributes, as individuals are often unwilling or unable to disclose such information due to privacy and ethical concerns (Dai and Wang, 2021). Therefore, addressing fairness under incomplete sensitive attributes is essential for real-world social network applications, where reliable and equitable outcomes must be guaranteed (Zhang et al., 2024b).
@@ 2.1. Graph Transformers
GTs have demonstrated strong potential in graph representation learning by providing global receptive fields (Shehzad et al., 2024; Lin et al., 2024; Zhang et al., 2024a). Models such as GraphTrans (Wu et al., 2021) and SAN (Kreuzer et al., 2021) strengthen long-range dependency modelling through attention mechanisms and spectral encodings. NAGphormer (Chen et al., 2023) extends this paradigm by treating nodes as sequences of token vectors and leveraging selected Laplacian eigenvectors to encode structural topology and enable multi-hop message aggregation. These developments have positioned Transformer-based graph models at the core of a wide range of social network applications requiring expressive, non-local reasoning.
@@ 2.2. Fairness-aware GNNs in Incomplete Social Networks
Fairness-aware GNNs have attracted growing attention, with methods designed to protect or regulate sensitive attributes during learning (Dong et al., 2022). Representative models like FairGB (Li et al., 2024) and FUGNN (Luo et al., 2024a) employ group balancing, counterfactual mixup, and spectral regularisation to jointly optimise fairness and utility. However, most of these approaches assume full access to sensitive attributes, which is often unrealistic in social network applications where such information is incomplete due to privacy constraints (Dai et al., 2023; Chen et al., 2024a).
To address this challenge, recent approaches such as FairGNN (Dai and Wang, 2023) and FairAC (Guo et al., 2023) integrate attribute estimation and debiasing strategies under incomplete sensitive attributes. However, these methods rely on generating or inferring sensitive attributes, which can reintroduce privacy risks by reconstructing deliberately withheld data. Furthermore, most fairness-aware designs are tightly coupled with GNN-specific operations, limiting their applicability to GTs, that encode structure differently and capture global dependencies (zhang2024; Xiang et al., 2025). Consequently, enabling fairness-aware GTs in incomplete social networks remains a critical yet underexplored challenge.
@@ 6.3. Comparison Results
To simulate incomplete sensitive features, this study uniformly masks values in the complete sensitive attribute column. This methodology is consistently applied across all models to ensure fair comparison, with both original and incomplete-data distributions explicitly defined in the experiments. Figure 3 and Figure 4 provides a detailed comparison of the fairness evaluation metrics between the proposed FairGE and several baseline models across seven social network datasets with varying missing rates. Specifically, comparison experiments are conducted with missing rates ranging from $10\%$ to $60\%$, covering both sparse and dense incompleteness scenarios. The overall accuracy, along with two key fairness metrics, $\Delta_{\text{SP}}$ and $\Delta_{\text{EO}}$, part of them are reported in Table 2, and others are shown in https://anonymous.4open.science/r/FairGE-609E.
The AMiner-S dataset contains multi-class sensitive attributes. While previous fairness evaluation metrics are formulated for binary scenarios, they can be readily extended to multi-class sensitive attribute setting, as demonstrated by previous work (Luo et al., 2025). To ensure fairness across all sensitive subgroups, we quantify the core principle using the variance between them. Specifically, the fairness evaluation metrics for multi-class sensitive attributes are defined as follows, where class number denotes $m$:

### v2

@@ 1. Introduction
Graph Transformers (GTs) address common Graph Neural Networks (GNNs) challenges such as over-smoothing (Zhou et al., 2025), over-squash-ing (Gravina et al., 2025), and long-range dependencies (Chen et al., 2024a) by employing global attention to capture long-range node interactions (Wu et al., 2024; Zhang et al., 2025). Despite these advantages, GTs often overlook inherent biases in graph data, leading to discriminatory predictions towards sensitive subgroups (e.g., gender and race) (Luo et al., 2024b). This issue is particularly critical in social networks, where fairness directly affects user trust and system reliability (Zhong et al., 2024; Bernard and Balog, 2025; Luo et al., 2025b). Here, algorithmic fairness requires that model predictions remain unbiased with respect to sensitive attributes (e.g., gender, race, region, and other potential factors), ensuring equitable treatment of diverse subgroups (Dong et al., 2023; Xu et al., 2022, 2025; Oldfield et al., 2025). The challenge is further compounded by the frequent absence of sensitive attributes, as individuals are often unwilling or unable to disclose such information due to privacy and ethical concerns (Dai and Wang, 2021; Xu et al., 2023). Therefore, addressing fairness under incomplete sensitive attributes is essential for real-world social network applications, where reliable and equitable outcomes must be guaranteed (Zhang et al., 2024c).
@@ 2.1. Graph Transformers
GTs have demonstrated strong potential in graph representation learning by providing global receptive fields (Shehzad et al., 2024; Lin et al., 2024; Zhang et al., 2024b). Models such as GraphTrans (Wu et al., 2021) and SAN (Kreuzer et al., 2021) strengthen long-range dependency modelling through attention mechanisms and spectral encodings. NAGphormer (Chen et al., 2023) extends this paradigm by treating nodes as sequences of token vectors and leveraging selected Laplacian eigenvectors to encode structural topology and enable multi-hop message aggregation. These advances have made Transformer-based graph models central to social network applications requiring expressive, non-local reasoning.
@@ 2.2. Fairness-aware GNNs in Incomplete Social Networks
Fairness-aware GNNs have attracted growing attention, with methods designed to protect or regulate sensitive attributes during learning (Dong et al., 2022). Representative models like FairGB (Li et al., 2024) and FUGNN (Luo et al., 2024a) employ group balancing, counterfactual mixup, and spectral regularisation to jointly optimise fairness and utility. However, most of these approaches assume full access to sensitive attributes, which is often unrealistic in social network applications where such information is incomplete due to privacy constraints (Dai et al., 2023; Chen et al., 2024b).
To address this challenge, recent approaches such as FairGNN (Dai and Wang, 2023) and FairAC (Guo et al., 2023) integrate attribute estimation and debiasing strategies under incomplete sensitive attributes. However, these methods rely on generating or inferring sensitive attributes, which can reintroduce privacy risks by reconstructing deliberately withheld data. Furthermore, most fairness-aware designs are tightly coupled with GNN-specific operations, limiting their applicability to GTs, that encode structure differently and capture global dependencies (Zhang et al., 2024d; Xiang et al., 2025; Zhang et al., 2024a). Consequently, enabling fairness-aware GTs in incomplete social networks remains a critical yet underexplored challenge.
@@ 6.3. Comparison Results
To simulate incomplete sensitive features, this study uniformly masks values in the complete sensitive attribute column. This methodology is consistently applied across all models to ensure fair comparison, with both original and incomplete-data distributions explicitly defined in the experiments. Figure 3 and Figure 4 provides a detailed comparison of the fairness evaluation metrics between the proposed FairGE and several baseline models across seven social network datasets with varying missing rates. Specifically, comparison experiments are conducted with missing rates ranging from $10\%$ to $60\%$, covering both sparse and dense incompleteness scenarios. The overall accuracy, along with two key fairness metrics, $\Delta_{\text{SP}}$ and $\Delta_{\text{EO}}$, part of them are reported in Table 2, and others are shown in https://github.com/LuoRenqiang/FairGE.
The AMiner-S dataset contains multi-class sensitive attributes. While previous fairness evaluation metrics are formulated for binary scenarios, they can be readily extended to multi-class sensitive attribute setting, as demonstrated by previous work (Luo et al., 2025a). To ensure fairness across all sensitive subgroups, we quantify the core principle using the variance between them. Specifically, the fairness evaluation metrics for multi-class sensitive attributes are defined as follows, where class number denotes $m$: